# LoopServe: An Adaptive Dual-phase LLM Inference Acceleration System for Multi-Turn Dialogues

### v1

@@ 6.3. Parameter Sensitivity
Figure 9. Parameter Sensitivity to $B$.
(a) Llama-3.1-8B-Instruct.
(b) Qwen-2.5-7B-Instruct.
Figure 10. Parameter Sensitivity to $n_{d}$
6.3.3. The decoding interval $n_{d}$ in Algorithm 3.
The decoding interval $n_{d}$ determines how frequently LoopServe re-selects important input tokens based on the most recent output during progressive KV compression. A smaller $n_{d}$ allows the system to adapt more frequently to evolving output dependencies, which can further boost accuracy in challenging, dynamic dialogues. However, this comes with increased overhead due to more frequent KV cache updates. Specifically, we conduct experiments on MultiNews and GovReport data with Llama and Qwen backbones, containing questions at the beginning (-B), middle (-M), and end (-E) positions. As shown in Figure 10, our sensitivity analysis finds that moderate values of $n_{d}$ (e.g., 16 or 32) provide the best balance: they preserve the majority of the efficiency gains while maintaining robust generation quality. Extremely large $n_{d}$ values reduce the adaptivity of the KV cache and may result in decreased relevance of selected tokens, leading to lower performance on complex, multi-turn tasks.
7. Conclusion

### v2

@@ A.8 Supplementary Experimental Figure for Parameter Sensitivity
(c) Llama-3.1-8B-Instruct.
(d) Qwen-2.5-7B-Instruct.
Figure 8: Parameter sensitivity to $\alpha$ in (a), budget $B$ in (b), and decoding interval $n_{d}$ in (c) and (d).
A.8.1 The decoding interval $n_{d}$ in Algorithm 3
The decoding interval $n_{d}$ controls how often LoopServe re-selects important input tokens during progressive KV compression. Smaller $n_{d}$ values enable frequent adaptation to changing output dependencies, improving accuracy in dynamic dialogues but increasing overhead from more KV cache updates. We evaluate this on MultiNews and GovReport with Llama-3.1-8B-Instruct. and Qwen-2.5-7B-Instruct., covering questions at the beginning, middle, and end. As shown in Figure 8 (c) and (d), moderate $n_{d}$ values (e.g., 16 or 32) strike the best balance, maintaining efficiency and robust generation quality. Very large $n_{d}$ reduces adaptivity, leading to lower performance on complex, multi-turn tasks.
A.9 The Usage of LLMs for Paper Writing
We use GPT-4o and DeepSeek-R1 to polish our paper.